# On stochastic forms of functional isoperimetric inequalities

### v1

@@ 1. Introduction
where $W_{i}(\cdot)$ denotes the $i$-th quermassintegral (see Section 2 for a precise definition) and $B_{K}$ is an Euclidean ball with the same volume as $K$. When $i=1$, (1.1) is nothing but the classical isoperimetric inequaliy for convex bodies. For context and background the reader may check the excellent books by R. Gardner [16], P. Gruber [19] and R. Schneider [41].
Inspired by the work of H. Groemer [18] on the expected volume of random polytopes, the isoperimetric-type inequality (1.1) was shown to hold, by M. Hartzoulaki and G. Paouris in [22], also in expectation when dealing with random convex sets. In a more detailed manner, for any collection of independent random vectors $\{X_{k}\}_{k=1}^{N}$ uniformly distributed on $K$, they proved that
@@
for every $i\in\{0,1,\dots,n-1\}$, where $\{X_{k}^{*}\}_{k=1}^{N}$ are independent random vectors uniformly distributed on $B_{K}$. Note that (1.2) recovers (1.1) as $N\to+\infty$.
A systematic study of stochastic dominance phenomena in isoperimetric-type inequalities was initiated in [31] (and further developed in [32, 13, 33, 1, 14]) where, among other findings, it was proven that convexity (or concavity) properties of certain functionals lead to an stochastic dominance in the associated inequalities. More recently, a similar program focused on stochastic analogues of functional inequalities in the realm $p$-concave functions was started by P. Pivovarov and J. Rebollo-Bueno in [36, 37]. The aim of this paper is to continue their program by presenting several stochastic analogues of functional isoperimetric inequalities for $p$-concave functions.
@@
In light of (1.2), it is natural to wonder about a stochastic analogue of (1.3). Our first main result provides such an analogue. To state it precisely, we first introduce the stochastic framework developed by P. Pivovarov and J. Rebollo-Bueno. Let $N\geq n+1$. Given an integrable $p$-concave function $f:\mathbb{R}^{n}\to\mathbb{R}_{+}$, with $p\in\mathbb{R}\cup\{+\infty\}$, and $\{(X_{k},Z_{k})\}_{k=1}^{N}$ independent random vectors uniformly distributed w.r.t. the Lebesgue measure on the hypograph of $f$, i.e., uniformly sampled on
@@
(for the sake of simplicity, in the following we will simply say distributed w.r.t. $f$). The so-called stochastic model of $f$ is build upon the random sets $E_{N,p},H_{N,p}\subset\mathbb{R}^{n}\times\mathbb{R}_{+}$ given by
@@
$R_{z}(x)=\{(x,r):r\geq z\}\quad\mathrm{and}\quad\widetilde{R}_{z}(x)=\{(x,z):r\leq z\}.$
@@
$\mathrm{epi}\bigl{(}(\Phi^{N}_{(X_{k},Z_{k})})^{p}\bigr{)}=E_{N,p}\quad\mathrm{if}\quad p<0,$
$\mathrm{epi}(-\log\Phi^{N}_{(X_{k},Z_{k})})=E_{N,p}\quad\mathrm{if}\quad p=0,$
$\mathrm{hyp}\bigl{(}(\Phi^{N}_{(X_{k},Z_{k})})^{p}\bigr{)}=H_{N,p}\quad\mathrm{if}\quad p>0.$
@@
$\mathbb{E}\left[W_{i}\Bigl{(}\Phi^{N}_{(X_{k},Z_{k})}\Bigr{)}\right]\geq\mathbb{E}\left[W_{i}\Bigl{(}\Phi^{N}_{(X^{*}_{k},Z^{*}_{k})}\Bigr{)}\right],$
@@
$\mathbb{E}\left[\mathrm{Per}\Bigl{(}\Phi^{N}_{(X_{k},Z_{k})}\Bigr{)}\right]\geq\mathbb{E}\left[\mathrm{Per}\Bigl{(}\Phi^{N}_{(X^{*}_{k},Z^{*}_{k})}\Bigr{)}\right],$
@@
To achieve this goal, we employ the notion of functional polar projection bodies, which originates from ideas of E. Lutwak, D. Yang and G. Zhang [27, 26] (see also [25, 21, 20]). This concept was also formally introduced in [2] following [23] (See Section 2 for a precise definition). Within this framework, Zhang’s affine Sobolev inequality takes the following form. Note that, as aforementioned, this inequality is known to hold without the $p$-concavity assumption.
@@
$\mathbb{E}\left[\nu\bigl{(}\Pi^{\circ}\Phi^{N}_{(X_{k},Z_{k})}\bigr{)}\right]\leq\mathbb{E}\left[\nu\bigl{(}\Pi^{\circ}\Phi^{N}_{(X^{*}_{k},Z^{*}_{k})}\bigr{)}\right],$
@@
We would like to remark that the proof of Theorem 1.2 follows the approach of [34], which itself builds upon the ideas presented in [28, Section 8.2]. For our purposes, we employ an interpretation of the functional polar projection bodies using the functional mixed volumes of V. Milman and L. Rotem [29]. Furthermore, both Theorem 1.1 and Theorem 1.2 recover their geometric counterparts established by M. Hartzoulaki and G. Paouris [22], and by G. Paouris, P. Pivovarov, and K. Tatarko [34] (see Remark 3.2), respectively. Given these connections, it is natural to ask whether Theorems A and B can also be recovered from their stochastic versions. We conclude by addressing this question.
@@
This paper is organized as follows. In Section 2 we collect some preliminaries, background material and several tools the we will use later on, while Section 3 is devoted to the proofs of Theorems 1.1 and 1.2. Finally, we show in Section 4 how to recover Theorems A and B from their stochastic versions.
@@ 2. Preliminaries
We work in the $n$-dimensional Euclidean space $\mathbb{R}^{n}$ endowed with the standard inner product $\left\langle\cdot,\cdot\right\rangle$. We write $e_{i}$ for the $i$-th canonical unit vector, and $x_{i}$ is used for the $i$-th coordinate of a vector in such a space. We denote by $B^{n}_{2}$ the $n$-dimensional Euclidean (closed) unit ball and by $\mathbb{S}^{n-1}$ its boundary. We use ++ for the Minkowski sum, i.e., $A+B=\{a+b:\,a\in A,\,b\in B\}$ for any non-empty sets $A,B\subset\mathbb{R}^{n}$. Given any set $M\subset\mathbb{R}^{n}$, $\chi_{{}_{M}}$ denotes its characteristic function. Moreover, for any unit direction $\in\mathbb{S}^{n-1}, uu^{\perp}$, $P_{u^{\perp}}$ and $R_{u}$ are used for a hiperplane with normal vector $u$, the orthogonal projection onto it, and the reflection map about it, respectively. Along the paper, the $k$-dimensional Lebesgue measure of a measurable set is denoted by $|\cdot|_{k}$ and we will omit the index $k$ when it is equal to the dimension of the ambient space; furthermore, as usual, integrating $\mathrm{d}x$ will stand the integration with respect to the Lebesgue measure.
Let $K\subset\mathbb{R}^{n}$ be a convex set, its support function $h_{K}$ and polar body $K^{\circ}$ are given by
@@
If $K$ contains the origin in its interior, its Minkowski functional is defined as $\|x\|_{K}=\inf\{\lambda>0:x\in\lambda K\}$. Note that, in this case, $K=\{x\in\mathbb{R}^{n}:\|x\|_{K}\leq 1\}$ and $\|\cdot\|_{K^{\circ}}=h_{K}(\cdot)$. We will also recall some definitions; the projection body, $\Pi K$, of $K$ is the centrally symmetric convex body whose support function is given by $h_{\Pi K}(u)=|P_{u^{\perp}}K|_{n-1}$ where, for any given direction $u\in\mathbb{S}^{n-1}$, $[0,u]$ is a segment joining the origin and $u$. Thus the polar projection body, denoted as $\Pi^{\circ}K$, is merely $(\Pi K)^{\circ}$. For convex sets $K,L\subset\mathbb{R}^{n}$ the Hausdorff distance between them, denoted as $\delta^{H}(K,L)$ is
$\delta^{H}(K,L)=\inf\{\varepsilon>0:M\subset L+\varepsilon B_{2}^{n},L\subset M+\varepsilon B_{2}^{n}\}.$
@@
where the coefficient $V(K_{i_{1}},...,K_{i_{n}})$ is the so-called mixed volume of $n$-tuple $(K_{i_{1}},...,K_{i_{n}})$ (see [41, Section 5] for background and properties). We will use the standard abbreviation V(K[nk],L[k])V(K[n-k],L[k]) when the sets appear, respectively, nkn-k and $k$ times in the mixed volume with $k\in\{1,...,n-1\}$. In addition, for any $i\in\{1,...,n-1\}$, the $i$-th quermassintegral of $K$, is denoted by $W_{i}(K)=V(K[n-i],B_{2}^{n}[i])$. Note that, with this terminology, $\|u\|_{\Pi^{\circ}K}=nV(K[n-1],[0,u])$ for every $u\in\mathbb{S}^{n-1}$. Moreover, we will use that mixed volumes are continuous w.r.t. the Hausdorff distance.
@@
It was proved in [17, Theorem 6.1] that, if $K_{1},\dots,K_{N}$ are convex and $M$ is a convex body contained in the positive orthant or centrally symmetric, then $\oplus_{M}(K_{1},\dots,K_{N})$ is convex.
We remark that the use of the $M$-addition allow us to recover certain random sets that have already been use in the literature when proving stochastic forms of isoperimetric inequalities (see e.g. [31] and [34]). In this regard, note that if $x_{1},\dots,x_{N}]$ is the $n\times N$ matrix with columns $x_{i}$ then
@@ 2.1. Linear parameter systems
We will use the notion of a linear parameter system introduced by C. A. Rogers and G. C. Shepard in [40] (see also [41, Section 10.4]). Let $K\subset\mathbb{R}^{n}$ be a convex body, and $u\in\mathbb{S}^{n-1}$ be a unit direction. A linear parameter system is a family of convex bodies $\{K(t)\}_{t\in I}$, where $I\subset\mathbb{R}$ is an interval, that can be represented as
@@
A fundamental property of linear parameter systems is their convexity under mixed volumes: an elegant proof can be found in [41, Theorem 10.4.1].
@@
Let $\{K(t)\}_{t\in I}$ a linear parameter system. Then for every $i=0,\dots,n-1$ $t\mapsto W_{i}\bigl{(}K(t)\bigr{)}$ is convex.
@@ 2.2. Functional setting
In the following we will assume that $p$-concave functions are upper semicontinuous. Indeed, otherwise we may replace the function by its upper closure, which is determined via the closure of the superlevel sets (see [39, Theorem 1.6]), and thus their Lebesgue measure is preserved. In particular, this will imply that the maximum of such a function is attained. Additionally, we will always consider integrable $p$-concave functions. The latter, together with upper semicontinuity, imply that in our framework the superlevel sets of a $p$-concave function are convex bodies.
@@
When $s=$-$\infty$ the measure is usually referred to as convex. Following Borell’s characterization [5], an absolutely continuous measure $\nu$ in $\mathbb{R}^{n}$ with density $\varphi$ is $s$-concave if and only if $\varphi$ is $p$-concave with p=s/(1ns)p=s/(1-ns) (Note that Jensen’s inequality implies that convex measures are the largest class among $s$-concave ones). The latter can be deduced from the following result, originally proved in [5] and [6] (see also [15] for a detailed presentation).
@@
We will also recall the notion of symmetric decreasing rearrangement. We essentially follow [8] (see also [24, Chapter 3]). Let $A\subset\mathbb{R}^{n}$ be a measurable set with finite volume. Its symmetric rearrangement $A^{*}$ is an Euclidean open ball with the same volume as $A$. Let now $f:\mathbb{R}^{n}\to\mathbb{R}_{+}$ be an integrable function. Using its layer-cake representation
@@
Note that, for any given $t\in\mathbb{R}_{+}$, $\{f\geq t\}$ is used to denote the superlevel set $\{x\in\mathbb{R}^{n}:f(x)\geq t\}$. We also emphasize that $f^{*}$ is radially symmetric and decreasing. In addition, it preserves the volume of the superlevel sets, i.e., $|\{f>t\}|=|\{f^{*}>t\}|$ for all $t\in\mathbb{R}_{+}$. Moreover, one has $\|f^{*}\|_{p}=\|f\|_{p}$ for each $1\leq p\leq+\infty$, where $\|\cdot\|_{p}$ is the usual $L_{p}(\mathbb{R}^{n})$-norm.
@@
Equivalently, $f^{u}$ is obtained rearranging $f$ along every line parallel to $u$, i.e., for every $y\in u^{\perp}$, taking $h(t)=f(y+tu)$, we have that $f^{u}(y+tu)=h^{*}(t)$. It is proved in [7] (see also [42, Chapter 14]) that for every measurable function $f:\mathbb{R}^{n}\to\mathbb{R}_{+}$ with compact support there exists a sequence of the form $f_{0}=f$ and $f_{n+1}=f_{n}^{u}$ for some $u\in\mathbb{S}^{n-1}$, which converges in the $L_{1}(\mathbb{R}^{n})$-norm to $f^{*}$.
@@
which implies (2.4) after a sequence of Steiner symmetrizations with respect to suitable directions. We refer the reader to [31, Proposition 3.2] for a detailed exposition of the latter. Note that, if $F_{u,y}$ is quasi-convex for any $u\in\mathbb{S}^{n-1}$ and $y=(y_{1},...,y_{N})\in u^{\perp}$, the inequalities (2.4) and (2.5) are reversed.

### v2

@@ 1. Introduction
where $W_{i}(\cdot)$ denotes the $i$-th quermassintegral (see Section 2 for a precise definition) and $B_{K}$ is a Euclidean ball with the same volume as $K$. When $i=1$, (1.1) is nothing but the classical isoperimetric inequality for convex bodies. For context and background the reader may check the excellent books by R. Gardner [17], P. Gruber [19] and R. Schneider [40].
Inspired by the work of H. Groemer [18] on the expected volume of random polytopes, the isoperimetric-type inequality (1.1) was shown to hold, by R. E. Pfiefer in [35] and by M. Hartzoulaki and G. Paouris in [22], in expectation when dealing with random convex sets (see also [22]). In a more detailed manner, for any collection of independent random vectors $\{X_{k}\}_{k=1}^{N}$ uniformly distributed in $K$, they proved that
@@
for every $i\in\{0,1,\dots,n-1\}$, where $\{X_{k}^{*}\}_{k=1}^{N}$ are independent random vectors uniformly distributed in $B_{K}$. Note that (1.2) recovers (1.1) as $N\to+\infty$.
A systematic study of stochastic dominance phenomena in isoperimetric-type inequalities was initiated in [30] (and further developed in [31, 13, 32, 1, 14]) where, among other findings, it was proven that convexity (or concavity) properties of certain functionals lead to a stochastic dominance in the associated inequalities. More recently, a similar program focused on stochastic analogues of functional inequalities in the realm of $p$-concave functions was started by P. Pivovarov and J. Rebollo-Bueno in [36, 37]. The aim of this paper is to continue their program by presenting several stochastic analogues of functional isoperimetric inequalities for $p$-concave functions.
@@
In light of (1.2), it is natural to wonder about a stochastic analogue of (1.3). Our first main result provides such an analogue. To state it precisely, we first introduce the stochastic framework developed by P. Pivovarov and J. Rebollo-Bueno. Let $N\geq n+1$. Given an integrable $p$-concave function $f:\mathbb{R}^{n}\to\mathbb{R}_{+}$, with $p\in\mathbb{R}\cup\{+\infty\}$, and $\{(X_{k},Z_{k})\}_{k=1}^{N}$ independent random vectors uniformly distributed w.r.t. the Lebesgue measure on the hypograph of $f$, i.e., uniformly sampled in
@@
(for the sake of simplicity, in the following we will simply say distributed w.r.t. $f$). The so-called stochastic model of $f$ is built upon the random sets $E_{N,p},H_{N,p}\subset\mathbb{R}^{n}\times\mathbb{R}_{+}$ given by
@@
$R_{z}(x)=\{(x,r):r\geq z\}\quad\mathrm{and}\quad\widetilde{R}_{z}(x)=\{(x,r):r\leq z\}.$
@@
(1.6) $\mathrm{epi}\bigl((\Phi^{N}_{(X_{k},Z_{k})})^{p}\bigr)=E_{N,p}\quad\mathrm{if}\quad p<0,$
(1.7) $\mathrm{epi}(-\log\Phi^{N}_{(X_{k},Z_{k})})=E_{N,p}\quad\mathrm{if}\quad p=0,$
(1.8) $\mathrm{hyp}\bigl((\Phi^{N}_{(X_{k},Z_{k})})^{p}\bigr)=H_{N,p}\quad\mathrm{if}\quad p>0.$
@@
[…]
@@
[…]
@@
To achieve this goal, we employ the notion of functional polar projection bodies, which originates from ideas of E. Lutwak, D. Yang and G. Zhang [26, 25] (see also [24, 21, 20]). This concept was also formally introduced in [10, 2] (see Section 2 for a precise definition). Within this framework, Zhang’s affine Sobolev inequality takes the following form. Note that, as mentioned, this inequality is known to hold without the $p$-concavity assumption.
@@
[…]
@@
We would like to remark that the proof of Theorem 1.2 follows the approach of [33], which itself builds upon the ideas presented in [27, Section 8.2]. For our purposes, we employ an interpretation of the functional polar projection bodies using the functional mixed volumes of V. Milman and L. Rotem [28]. Furthermore, Theorem 1.1 recovers the geometric counterparts established by R. E. Pfiefer [35] and by M. Hartzoulaki and G. Paouris [22], while Theorem 1.2 recovers that of G. Paouris, P. Pivovarov, and K. Tatarko [33] (see Remark 3.2).
Given these connections, it is natural to ask whether Theorems A and B can also be recovered from their stochastic versions. We conclude by addressing this question, noting first that the stochastic models are well-defined for integrable functions, without requiring any concavity assumption. However, the resulting random sets (see (1.6)) need not approximate the original function unless it is $p$-concave. Hence, the $p$-concavity hypothesis in Theorems 1.1 and 1.2 is imposed to ensure that the deterministic inequalities can be recovered from their stochastic counterparts.
@@
This paper is organized as follows. In Section 2 we collect some preliminaries, background material and several tools that we will use later on, while Section 3 is devoted to the proofs of Theorems 1.1 and 1.2. Finally, we show in Section 4 how to recover Theorems A and B from their stochastic versions.
@@ 2. Preliminaries
We work in the $n$-dimensional Euclidean space $\mathbb{R}^{n}$ endowed with the standard inner product $\left\langle\cdot,\cdot\right\rangle$. We write $e_{i}$ for the $i$-th canonical unit vector, and $x_{i}$ is used for the $i$-th coordinate of a vector in such a space. We denote by $B^{n}_{2}$ the $n$-dimensional Euclidean (closed) unit ball and by $\mathbb{S}^{n-1}$ its boundary. We use ++ for the Minkowski sum, i.e., $A+B=\{a+b:\,a\in A,\,b\in B\}$ for any non-empty sets $A,B\subset\mathbb{R}^{n}$. Given any set $M\subset\mathbb{R}^{n}$, $\chi_{{}_{M}}$ denotes its characteristic function. Moreover, for any unit direction $\in\mathbb{S}^{n-1}, uu^{\perp}$, $P_{u^{\perp}}$ and $R_{u}$ are used for a hyperplane with normal vector $u$, the orthogonal projection onto it, and the reflection map about it, respectively. Along the paper, the $k$-dimensional Lebesgue measure of a measurable set is denoted by $|\cdot|_{k}$ and we will omit the index $k$ when it is equal to the dimension of the ambient space; furthermore, as usual, integrating $\mathrm{d}x$ will stand the integration with respect to the Lebesgue measure.
Let $K\subset\mathbb{R}^{n}$ be a compact convex set, its support function $h_{K}$ and polar body $K^{\circ}$ are given by
@@
If $K$ contains the origin in its interior, its Minkowski functional is defined as $\|x\|_{K}=\inf\{\lambda>0:x\in\lambda K\}$. Note that, in this case, $K=\{x\in\mathbb{R}^{n}:\|x\|_{K}\leq 1\}$ and $\|\cdot\|_{K^{\circ}}=h_{K}(\cdot)$. We will also recall some definitions; the projection body, $\Pi K$, of $K$ is the centrally symmetric convex body whose support function is given by $h_{\Pi K}(u)=|P_{u^{\perp}}K|_{n-1}$. Thus the polar projection body, denoted as $\Pi^{\circ}K$, is merely $(\Pi K)^{\circ}$. For convex sets $K,L\subset\mathbb{R}^{n}$ the Hausdorff distance between them, denoted as $\delta^{H}(K,L)$ is
$\delta^{H}(K,L)=\inf\{\varepsilon>0:K\subset L+\varepsilon B_{2}^{n},L\subset K+\varepsilon B_{2}^{n}\}.$
@@
where the coefficient $V(K_{i_{1}},...,K_{i_{n}})$ is the so-called mixed volume of $n$-tuple $(K_{i_{1}},...,K_{i_{n}})$ (see [40, Section 5] for background and properties). We will use the standard abbreviation V(K[nk],L[k])V(K[n-k],L[k]) when the sets appear, respectively, nkn-k and $k$ times in the mixed volume with $k\in\{1,...,n-1\}$. In addition, for any $i\in\{1,...,n-1\}$, the $i$-th quermassintegral of $K$, is denoted by $W_{i}(K)=V(K[n-i],B_{2}^{n}[i])$. Note that, with this terminology, $\|u\|_{\Pi^{\circ}K}=nV(K[n-1],[0,u])$ for every $u\in\mathbb{S}^{n-1}$. Moreover, we will use that mixed volumes are continuous w.r.t. the Hausdorff distance.
@@
It was proved in [16, Theorem 6.1] that, if $K_{1},\dots,K_{N}$ are convex and $M$ is a compact convex set contained in the positive orthant or centrally symmetric, then $\oplus_{M}(K_{1},\dots,K_{N})$ is convex.
We remark that the use of the $M$-addition allows us to recover certain random sets that have already been used in the literature when proving stochastic forms of isoperimetric inequalities (see e.g. [30] and [33]). In this regard, note that if $x_{1},\dots,x_{N}]$ is the $n\times N$ matrix with columns $x_{i}$ then
@@ 2.1. Linear parameter systems
We will use the notion of a linear parameter system introduced by C. A. Rogers and G. C. Shepard in [39] (see also [40, Section 10.4]). Let $K\subset\mathbb{R}^{n}$ be a convex body, and $u\in\mathbb{S}^{n-1}$ be a unit direction. A linear parameter system is a family of convex bodies $\{K(t)\}_{t\in I}$, where $I\subset\mathbb{R}$ is an interval, that can be represented as
@@
A fundamental property of linear parameter systems is their convexity under mixed volumes: an elegant proof can be found in [40, Theorem 10.4.1].
@@
Let $\{K(t)\}_{t\in I}$ be a linear parameter system. Then for every $i=0,\dots,n-1$ $t\mapsto W_{i}\bigl(K(t)\bigr)$ is convex.
@@ 2.2. Functional setting
In the following we will assume that $p$-concave functions are upper semicontinuous. Indeed, otherwise we may replace the function by its upper closure, which is determined via the closure of the superlevel sets (see [38, Theorem 1.6]), and thus their Lebesgue measure is preserved. In particular, this will imply that the maximum of such a function is attained. Additionally, we will always consider integrable $p$-concave functions. The latter, together with upper semicontinuity, imply that in our framework the superlevel sets of a $p$-concave function are convex bodies.
@@
When $s=$-$\infty$, the measure is usually referred to as convex. Following Borell’s characterization [5], an absolutely continuous measure $\nu$ in $\mathbb{R}^{n}$ with density $\varphi$ is $s$-concave if and only if $\varphi$ is $p$-concave with p=s/(1ns)p=s/(1-ns) (Note that Jensen’s inequality implies that convex measures are the largest class among $s$-concave ones). The latter can be deduced from the following result, originally proved in [5] and [6] (see also [15] for a detailed presentation).
@@
We will also recall the notion of symmetric decreasing rearrangement. We essentially follow [8] (see also [23, Chapter 3]). Let $A\subset\mathbb{R}^{n}$ be a measurable set with finite volume. Its symmetric rearrangement $A^{*}$ is a Euclidean open ball with the same volume as $A$. Let now $f:\mathbb{R}^{n}\to\mathbb{R}_{+}$ be an integrable function. Using its layer-cake representation
@@
Note that, for any given $t\in\mathbb{R}_{+}$, $\{f\geq t\}$ is used to denote the superlevel set $\{x\in\mathbb{R}^{n}:f(x)\geq t\}$. We also emphasize that $f^{*}$ is radially symmetric and decreasing. In addition, it preserves the volume of the superlevel sets, i.e., $|\{f>t\}|=|\{f^{*}>t\}|$ for all $t\in\mathbb{R}_{+}$. Moreover, one has $\|f^{*}\|_{p}=\|f\|_{p}$ for each $1\leq p\leq+\infty$, where $\|\cdot\|_{p}$ is the usual $L_{p}(\mathbb{R}^{n})$-norm. It is also worth mentioning that the symmetric decreasing rearrangement preserves log-concavity (see [28, Proposition 24]).
@@
Equivalently, $f^{u}$ is obtained rearranging $f$ along every line parallel to $u$, i.e., for every $y\in u^{\perp}$, taking $h(t)=f(y+tu)$, we have that $f^{u}(y+tu)=h^{*}(t)$. It is proved in [7] (see also [41, Chapter 14]) that for every measurable function $f:\mathbb{R}^{n}\to\mathbb{R}_{+}$ with compact support there exists a sequence of the form $f_{0}=f$ and $f_{n+1}=f_{n}^{u}$ for some $u\in\mathbb{S}^{n-1}$, which converges in the $L_{1}(\mathbb{R}^{n})$-norm to $f^{*}$.
@@
which implies (2.4) after a sequence of Steiner symmetrizations with respect to suitable directions. We refer the reader to [30, Proposition 3.2] for a detailed exposition of the latter. Note that, if $F_{u,y}$ is quasi-convex for any $u\in\mathbb{S}^{n-1}$ and $y=(y_{1},...,y_{N})\in(u^{\perp})^{N}$, the inequalities (2.4) and (2.5) are reversed.